Given the following information

C= 100+0.6 Yd

I = 500

G = 800

T = 200

X = 200

M = 50 + 0.3 Y

Find :

1. Equilibrium level of GDP in the economy

2. Marginal propensity to save in the economy

3. What happens to the equilibrium output when autonomous investment increases by 100 rupees? What is the value of the investment multiplier?


1
Expert's answer
2022-04-05T09:55:18-0400

1). Equilibrium output, Y, in an economy is found simply by:

"Y=C+I+G+(X-M)"

Given

C= 100+0.6 Yd

I = 500

G = 800

T = 200

X = 200

M = 50 + 0.3 Y

"C=100+0.6(Y-200)"

"C=-20+0.6Y"

"Y=-20+0.6Y+500+800+200-50-0.3Y"

"Y-0.6Y+0.3Y=-20+500+800+200-50"

"0.7Y=1430"

"Y=2042.85"


2) "MPS=1-MPC"

MPC=0.6

"MPS=1-0.6=0.4"


3) When autonomous investment increases by 100 rupees, output would also increase by 250 rupees. The multiplier is

"m=\frac{1}{MPS}"

"m=\frac{1}{0.4}"

"m=2.5"

"\Delta{Y}=m\Delta{I}"

"\Delta{Y}=2.5×100"

"\Delta{Y}=250"

The new equilibrium output becomes 2042.85+250=2292.85